OpenStudy (anonymous):

The volume of a cone that fits exactly inside a cylinder is 15 cubic feet. What is the volume of the cylinder? (5 points) 45 cubic feet 75 cubic feet 225 cubic feet 1500 cubic feet

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Compare the two volume formulas: \[V _{cylinder}=\pi r^2h\] \[V _{cone}=\frac{1}{3}\pi r^2h\]

OpenStudy (anonymous):

I don't know how to find the radius or the height

OpenStudy (mertsj):

If the volume of the cone is exactly 1/3 the volume of the cylinder, what should you do to the volume of the cone to get the volume of the cylinder?

OpenStudy (anonymous):

Multiply it by 3?

OpenStudy (mertsj):

Very good!!!

OpenStudy (mertsj):

Do that and you will have your answer.

OpenStudy (anonymous):

So multiply 15 by 3?

OpenStudy (mertsj):

That's what you said.

OpenStudy (anonymous):

ok so it would be 45 then, right?

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

ok thx that was easier than I thought.

OpenStudy (mertsj):

yw